OpenStudy (anonymous):

Say Frank can build a model airplane in 3 days, Jo can do it in 2, but Moe can only do it in 5 days. If they work together, they’ll build a model plane in: Possible Answers Selected Possible Answer 0.5 days 0.76 days 10 days 0.97 days 2 days

OpenStudy (anonymous):

i think its 0.76 im not sure really i think its the rule of 3

OpenStudy (anonymous):

a/b = c/x

OpenStudy (anonymous):

nope its not 0.76 thats what i guessed.

OpenStudy (anonymous):

would it be 2 days? since that is the fastest jo can do it

OpenStudy (anonymous):

thats the best i can think of sorry

OpenStudy (anonymous):

\[T_{F} = 3, T_{Jo} = 2, T_{Moe} = 5\]\[R_{F} = W/T_{F}, R_{Jo} = W/T_{Jo}, R_{M} = W/T_{M}\] W is work to be done to build a model airplane, R rate of work of wach person when three work together time taken \[W/(W/T_{F} +W/T_{Jo} +W/T_{M})\] 0.97 days

OpenStudy (anonymous):

@sumanth4phy what formula is that? Rate = work / time? could you explain what you did there?

OpenStudy (anonymous):

I see how you add their separate rates but why what are you dividing?

OpenStudy (anonymous):

yeah Rate/Speed at which work is done = W/T I need time so T=W/R hence the division

OpenStudy (anonymous):

OH gotcha riiight